OpenStudy (anonymous):

How much longer than 5/8 inch is 7/8?

OpenStudy (anonymous):

2/8

OpenStudy (anonymous):

It was wrong.

OpenStudy (anonymous):

The question has no restriction regarding simplification. 2/8 or 1/4 is correct.

OpenStudy (mathmath333):

\(\Large \begin{align} \color{black}{\dfrac{7}{8}-\dfrac{5}{8}\\~\\ =\dfrac{7-5}{8}=\dfrac{2}{8}=\dfrac{1}{4}} \end{align}\)